# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

### v1

@@ F.2 Cultural probing results
Table 6: Probing performance comparison with Arabic prompts and FmLAMA-ar sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural group within the same model. The Pearson correlation coefficient between mAP and mWS is 0.71 based on the results above.
Origin Count mB-c mB-u XRb XRl mT5 Avg.
Italy 59 (12.7%) 5.73±2.00 3.62±1.43 3.60±0.16 4.04±0.85 4.72±0.60 4.34
U.S. 67 (14.5%) 11.59±2.64 6.72±2.59 4.51±0.81 5.05±0.58 11.55±2.31 7.88
Turkey 12 (2.6%) 9.29±7.33 9.31±5.64 8.83±0.45 9.00±0.78 13.64±7.43 10.01
Japan 19 (4.1%) 8.37±4.61 5.81±3.26 7.39±1.97 8.02±0.77 8.56±3.44 7.63
France 65 (14.0%) 9.26±1.15 3.55±1.25 3.61±0.25 4.03±0.54 4.63±1.26 5.02
U.K. 23 (5.0%) 16.17±3.36 9.89±4.31 3.54±0.78 3.43±0.84 8.02±1.97 8.21
Mexico 11 (2.4%) 2.38±0.57 2.30±0.86 4.90±1.97 4.93±1.48 2.59±1.09 3.42
India 18 (3.9%) 4.83±3.24 5.68±6.31 5.30±2.41 6.45±0.15 6.30±4.71 5.71
Germany 11 (2.4%) 15.41±5.12 8.46±4.91 2.97±0.33 3.67±1.64 14.20±4.70 8.94
China 9 (1.9%) 3.94±3.35 2.63±1.49 6.12±1.36 6.47±0.97 4.19±4.65 4.67
Iran 6 (1.3%) 16.57±7.01 7.54±6.06 6.15±2.83 6.27±2.68 6.49±4.44 8.60
Greece 6 (1.3%) 4.33±2.31 1.85±0.60 6.73±2.89 6.82±2.87 4.24±1.18 4.79
Spain 7 (1.5%) 5.17±2.75 2.74±0.72 5.54±2.15 5.77±1.56 7.62±3.32 5.37
Russia 4 (0.9%) 5.22±2.97 4.82±3.18 5.91±0.27 5.44±0.84 3.53±1.26 4.98
Others 146 (31.5%) 5.48±0.57 3.60±1.46 4.94±0.84 5.15±0.43 4.38±1.34 4.71
ALL 463 (100.0%) 7.90±1.15 4.77±1.90 4.70±0.73 5.05±0.08 6.42±1.33 5.77
@@
mB-c mB-u XRb XRl mT5 Average
Italy 0.3313±0.01 0.3286±0.03 0.2963±0.00 0.3032±0.02 0.2987±0.07 0.3116
U.S. 0.3706±0.02 0.3442±0.04 0.3090±0.01 0.3130±0.02 0.3565±0.07 0.3387
Turkey 0.3085±0.07 0.3182±0.05 0.3048±0.04 0.2958±0.01 0.2868±0.15 0.3028
Japan 0.2963±0.04 0.2998±0.06 0.2765±0.02 0.2708±0.01 0.2868±0.09 0.2860
France 0.3429±0.01 0.3278±0.04 0.2987±0.02 0.2948±0.01 0.3089±0.05 0.3146
U.K. 0.3832±0.04 0.3776±0.02 0.2777±0.03 0.2681±0.01 0.3217±0.06 0.3257
Mexico 0.3585±0.02 0.3534±0.09 0.3444±0.03 0.3260±0.01 0.3605±0.06 0.3486
India 0.3401±0.03 0.3431±0.06 0.2809±0.02 0.2901±0.04 0.2738±0.13 0.3056
Germany 0.3703±0.06 0.3536±0.04 0.2923±0.06 0.2829±0.04 0.3623±0.07 0.3323
China 0.2713±0.02 0.3245±0.05 0.2646±0.04 0.2531±0.02 0.2541±0.12 0.2735
Iran 0.3856±0.08 0.3494±0.08 0.3827±0.00 0.3642±0.04 0.3724±0.06 0.3709
Greece 0.3310±0.02 0.3040±0.02 0.3393±0.02 0.3388±0.02 0.3286±0.05 0.3283
Spain 0.3128±0.03 0.3768±0.06 0.3646±0.01 0.3488±0.03 0.4115±0.10 0.3629
Russia 0.3431±0.04 0.3713±0.07 0.3803±0.04 0.3596±0.00 0.3268±0.06 0.3562
Others 0.3048±0.01 0.3105±0.05 0.2890±0.02 0.2842±0.01 0.2782±0.08 0.2933
ALL 0.3321±0.01 0.3287±0.04 0.2979±0.02 0.2952±0.01 0.3068±0.07 0.3121
@@
Table 7: Probing performance comparison with Hebrew prompts and FmLAMA-he sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural group within the same model. The Pearson correlation coefficient between mAP and mWS is 0.31 based on the results above.
Origin Count Bb-ru mB-c mB-u XRb XRl mT5 Avg.
Italy 101 (10.5%) 2.97±2.52 6.28±3.53 8.19±3.72 1.79±0.13 1.61±0.29 4.35±1.19 4.20
U.S. 79 (8.2%) 6.86±2.93 8.24±3.97 11.05±3.62 2.29±0.82 1.81±0.15 5.49±1.63 5.96
Turkey 28 (2.9%) 4.12±3.32 4.19±2.62 7.92±0.85 1.98±0.06 1.82±0.50 3.82±3.07 3.98
Japan 68 (7.1%) 4.41±2.32 3.79±1.22 5.37±2.72 1.67±0.22 1.52±0.02 1.91±1.15 3.11
France 93 (9.7%) 4.36±2.40 3.92±2.44 4.82±3.74 1.91±0.67 1.52±0.16 4.86±3.00 3.56
U.K. 33 (3.4%) 6.03±6.40 7.81±3.94 9.74±3.14 3.14±3.90 1.37±0.19 9.79±6.82 6.31
Mexico 22 (2.3%) 3.84±1.45 1.26±0.72 1.95±0.93 1.30±0.51 1.02±0.08 2.84±0.99 2.04
India 21 (2.2%) 5.86±5.73 9.58±5.82 15.71±6.27 1.55±0.43 1.49±0.55 5.29±4.03 6.58
Germany 45 (4.7%) 4.84±4.05 6.92±2.61 9.73±3.90 2.20±1.63 1.38±0.17 5.09±2.00 5.03
China 30 (3.1%) 10.09±6.49 6.93±2.75 10.59±5.73 1.29±0.99 0.83±0.09 5.08±3.75 5.80
Iran 13 (1.4%) 8.25±5.29 6.80±5.72 8.39±7.91 1.53±0.56 1.39±0.53 7.92±6.33 5.71
Greece 18 (1.9%) 3.54±4.18 0.99±0.22 1.15±0.43 1.48±0.09 1.51±0.14 4.09±1.45 2.13
Spain 38 (4.0%) 2.50±1.31 4.83±2.29 6.17±2.48 1.25±0.17 1.25±0.04 4.64±1.45 3.44
Russia 45 (4.7%) 3.61±1.79 3.23±1.01 4.10±2.41 2.18±0.89 2.16±0.69 2.90±0.74 3.03
Others 327 (34.0%) 3.36±2.13 4.31±1.80 5.79±1.31 1.77±0.15 1.57±0.29 3.68±1.00 3.41
ALL 961 (100.0%) 4.29±2.63 5.07±1.92 6.89±1.46 1.85±0.49 1.55±0.16 4.29±1.62 3.99
@@
Bb-ru mB-c mB-u XRb XRl mT5 Average
Italy 0.3229±0.06 0.3243±0.06 0.3730±0.06 0.2217±0.05 0.2206±0.05 0.3660±0.03 0.3048
U.S. 0.3555±0.05 0.3370±0.07 0.3895±0.03 0.2439±0.06 0.2336±0.03 0.3844±0.03 0.3240
Turkey 0.3129±0.05 0.3006±0.06 0.3554±0.05 0.2018±0.07 0.1915±0.04 0.3472±0.05 0.2849
Japan 0.2609±0.04 0.2622±0.02 0.2906±0.01 0.1619±0.04 0.1618±0.04 0.2712±0.01 0.2348
France 0.3233±0.05 0.2967±0.04 0.3378±0.03 0.2126±0.05 0.2099±0.04 0.3825±0.05 0.2938
U.K. 0.3513±0.08 0.3186±0.07 0.3735±0.03 0.2208±0.08 0.2059±0.05 0.4273±0.07 0.3162
Mexico 0.3264±0.03 0.3010±0.04 0.3348±0.04 0.2480±0.05 0.2431±0.03 0.3815±0.06 0.3058
India 0.2936±0.07 0.3112±0.09 0.4053±0.04 0.1877±0.05 0.1862±0.05 0.3235±0.04 0.2846
Germany 0.3403±0.06 0.3267±0.06 0.3825±0.07 0.2064±0.07 0.1894±0.03 0.3722±0.04 0.3029
China 0.3415±0.08 0.3144±0.04 0.3596±0.04 0.1873±0.06 0.1842±0.05 0.3515±0.04 0.2898
Iran 0.3358±0.07 0.3156±0.05 0.3699±0.06 0.2099±0.06 0.2020±0.04 0.4037±0.05 0.3061
Greece 0.3058±0.11 0.2960±0.05 0.3348±0.02 0.2252±0.04 0.2303±0.05 0.3716±0.02 0.2939
Spain 0.3287±0.04 0.3035±0.06 0.3532±0.03 0.2425±0.04 0.2406±0.03 0.3946±0.04 0.3105
Russia 0.3495±0.03 0.3152±0.05 0.3489±0.04 0.2590±0.05 0.2575±0.05 0.4177±0.03 0.3246
Others 0.3240±0.05 0.3105±0.05 0.3489±0.04 0.2186±0.05 0.2162±0.04 0.3660±0.03 0.2974
ALL 0.3245±0.05 0.3098±0.05 0.3536±0.03 0.2173±0.05 0.2132±0.04 0.3674±0.03 0.2976

### v3

@@ D.1 Probing results in each FmLAMA-la
Table 9: Probing performance comparison with Arabic prompts and FmLAMA-ar sub-dataset.
Origin Count Bb-zh mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 49 (8.7%) 17.18±1.67 13.07±4.92 10.55±0.84 24.54±2.87 9.32±4.48 13.58±4.08 14.71
U.S. 101 (17.9%) 19.32±1.47 13.79±3.56 12.37±1.09 26.50±2.94 12.91±3.61 13.17±6.45 16.34
Turkey 12 (2.1%) 8.92±3.86 6.30±2.12 5.22±0.18 18.16±3.49 8.18±3.58 11.00±6.39 9.63
Japan 114 (20.2%) 15.79±2.32 11.18±4.41 7.85±0.60 17.67±2.73 11.15±2.66 11.78±4.54 12.57
France 70 (12.4%) 19.45±4.13 16.04±5.31 9.14±0.98 20.96±1.45 13.36±3.62 14.71±5.89 15.61
U.K. 38 (6.7%) 27.88±2.65 19.55±7.36 10.64±0.68 28.43±4.16 18.02±8.33 21.46±6.75 21.00
Mexico 19 (3.4%) 21.12±0.86 13.40±3.51 12.05±0.79 16.25±4.70 10.60±1.32 8.21±3.83 13.61
India 24 (4.3%) 5.47±1.65 5.02±2.68 1.13±0.17 12.37±3.79 4.57±2.20 6.58±5.13 5.86
Germany 16 (2.8%) 10.31±2.53 12.46±3.54 6.81±1.75 10.41±2.12 4.08±2.38 7.62±4.63 8.62
China 87 (15.4%) 22.55±4.86 18.32±6.75 7.88±0.91 33.48±6.15 13.83±3.72 19.96±9.42 19.34
Iran 7 (1.2%) 7.72±1.44 8.19±4.58 3.54±0.20 8.39±2.08 3.40±0.38 8.82±3.50 6.68
Greece 7 (1.2%) 4.11±0.68 2.88±1.08 5.43±0.43 14.81±3.05 1.40±0.37 3.46±1.23 5.35
Spain 12 (2.1%) 13.00±1.83 10.21±1.58 12.72±1.17 21.02±0.93 5.53±3.77 12.41±6.05 12.48
Russia 8 (1.4%) 5.58±0.93 3.98±0.76 2.29±0.38 5.89±1.09 4.06±2.84 5.14±1.07 4.49
ALL 564 (100%) 17.84±2.37 13.56±4.36 8.96±0.44 22.71±2.37 11.46±3.32 13.76±5.56 14.71
Table 10: Probing performance comparison with Chinese prompts and FmLAMA-zh sub-dataset.
Origin Count mB mT5 Qwen2 Llama2 Llama3 Average
Italy 59 (18.6%) 4.18±2.54 4.30±1.49 7.78±2.37 3.97±1.47 6.29±3.14 5.30
U.S. 67 (21.1%) 8.62±7.22 8.13±2.02 9.09±4.52 3.62±0.81 10.17±4.44 7.93
Turkey 12 (3.9%) 8.99±8.65 6.33±4.01 9.21±5.92 2.93±1.18 10.68±6.33 7.63
Japan 19 (6.0%) 6.69±5.23 5.99±3.64 6.60±4.19 6.92±3.85 4.39±3.12 6.12
France 65 (20.5%) 5.43±3.01 4.54±0.92 6.40±2.74 2.72±0.54 5.50±1.47 4.92
U.K. 23 (7.3%) 9.43±6.79 6.95±2.26 9.79±5.06 3.97±2.49 10.08±5.66 8.04
Mexico 11 (3.5%) 2.22±0.75 2.26±0.74 4.33±2.44 2.87±1.99 2.14±2.01 2.76
India 18 (5.7%) 8.12±3.72 3.31±1.07 2.80±2.77 3.26±1.95 4.73±2.46 4.44
Germany 11 (3.5%) 7.57±5.98 14.84±6.23 8.59±3.46 5.92±1.83 9.38±5.42 9.26
China 9 (2.8%) 3.61±3.34 2.12±0.78 3.75±3.62 1.64±1.29 4.05±1.62 3.03
Iran 6 (1.9%) 8.58±5.37 5.05±2.90 19.40±9.90 2.95±0.57 20.63±10.96 11.32
Greece 6 (1.9%) 2.76±0.60 4.31±0.49 12.96±9.55 7.67±2.85 3.34±2.19 6.21
Spain 7 (2.2%) 4.58±1.81 7.02±1.87 3.86±1.66 4.04±1.62 2.82±1.26 4.46
Russia 4 (1.3%) 4.31±3.65 3.35±0.07 2.10±0.28 2.56±0.46 2.56±0.44 2.98
ALL 317 (100%) 6.41±4.30 5.77±1.59 7.57±2.66 3.74±0.85 7.17±2.63 6.13
@@
Origin mB mT5 Qwen2 Llama2 Llama3 Average
Italy 0.3102±0.03 0.3320±0.03 0.3197±0.03 0.2973±0.04 0.3265±0.05 0.3171
U.S. 0.3435±0.06 0.3626±0.03 0.3524±0.03 0.3042±0.03 0.3517±0.04 0.3429
Turkey 0.2976±0.06 0.3225±0.05 0.3110±0.04 0.2621±0.05 0.3296±0.10 0.3046
Japan 0.3134±0.05 0.3167±0.04 0.3114±0.05 0.2839±0.03 0.2899±0.02 0.3031
France 0.3082±0.03 0.3509±0.04 0.3334±0.02 0.2838±0.04 0.3134±0.03 0.3179
U.K. 0.3400±0.06 0.3418±0.04 0.3414±0.04 0.2930±0.02 0.3167±0.06 0.3266
Mexico 0.2959±0.05 {0.3829±0.06} 0.3420±0.06 0.2847±0.06 0.2865±0.04 0.3184
India {0.3614±0.05} 0.3437±0.06 0.3354±0.01 0.3133±0.04 0.3192±0.05 0.3346
Germany {0.3595±0.04} 0.3763±0.02 {0.3756±0.03} {0.3339±0.05} {0.3586±0.06} 0.3608
China 0.2863±0.03 0.3131±0.06 0.2714±0.06 0.2683±0.03 0.2432±0.04 0.2765
Iran 0.3435±0.05 0.3521±0.07 {0.4569±0.07} 0.2902±0.02 {0.4416±0.09} 0.3769
Greece 0.3014±0.02 0.3423±0.02 0.3722±0.08 0.3332±0.03 0.2951±0.04 0.3288
Spain 0.3493±0.03 {0.4229±0.07} 0.3547±0.02 {0.3635±0.06} 0.3260±0.03 0.3633
Russia 0.3276±0.03 0.3748±0.06 0.3319±0.04 0.3308±0.01 0.3179±0.04 0.3366
ALL 0.3237±0.04 0.3484±0.03 0.3369±0.02 0.2969±0.03 0.3248±0.04 0.3261
@@
Table 11: Probing performance comparison with Hebrew prompts and FmLAMA-he sub-dataset.
Origin Count Bb-ky Bb-kl mB-u mT5 Qwen2 Llama2 Llama3 Avg.
Italy 88 (16.7%) 6.90±1.57 7.75±2.34 3.85±2.54 6.02±0.64 3.50±1.10 1.19±0.22 2.23±0.75 4.49
U.S. 64 (12.2%) 11.37±2.66 10.87±3.86 6.32±5.20 10.15±0.34 2.46±0.69 1.68±0.12 2.95±0.93 6.54
Turkey 19 (3.6%) 2.75±2.47 4.13±2.62 4.07±3.17 1.82±0.44 3.86±1.00 2.16±0.12 7.47±1.45 3.75
Japan 101 (19.2%) 7.34±2.12 7.74±1.86 1.93±1.24 5.25±0.99 6.56±5.34 2.19±0.22 2.62±1.15 4.80
France 67 (12.7%) 2.33±0.43 3.09±0.44 3.15±0.17 2.79±0.30 3.10±0.68 1.40±0.09 2.41±1.27 2.61
U.K. 27 (5.1%) 14.07±2.83 15.89±4.43 6.00±4.42 7.55±1.34 3.21±1.02 1.79±0.31 6.89±5.70 7.91
Mexico 13 (2.5%) 5.18±3.69 8.37±2.85 4.89±3.29 2.31±0.40 7.12±3.13 2.75±1.99 4.18±1.61 4.97
India 32 (6.1%) 7.98±2.29 5.74±1.32 3.28±2.09 2.96±0.61 7.69±6.42 2.01±0.32 2.36±0.67 4.57
Germany 15 (2.9%) 2.50±0.94 5.08±1.25 2.42±1.68 3.20±1.18 2.63±0.95 1.33±0.48 1.91±0.21 2.72
China 54 (10.3%) 9.02±2.76 10.12±3.04 2.36±2.22 4.05±1.50 4.43±1.74 2.16±0.48 3.06±1.48 5.03
Iran 8 (1.5%) 7.99±5.23 3.76±1.41 1.62±0.67 1.52±0.24 1.50±0.14 2.87±2.49 2.12±1.83 3.05
Greece 9 (1.7%) 2.32±0.73 6.76±3.73 2.74±0.76 5.76±4.22 4.59±1.31 3.67±2.94 3.35±1.87 4.17
Spain 21 (4.0%) 4.05±2.25 5.90±2.70 1.79±1.24 5.82±1.33 3.99±1.38 2.50±0.97 3.85±0.99 3.99
Russia 8 (1.5%) 3.70±0.99 4.22±1.83 2.68±1.12 2.12±0.05 3.01±1.41 1.72±0.20 3.07±1.46 2.93
ALL 526 (100%) 7.05±1.47 7.68±2.00 3.45±2.10 5.19±0.42 4.32±1.76 1.86±0.18 3.07±0.95 4.66
Table 12: Probing performance comparison on mAP (%) with Korean prompts and FmLAMA-ko sub-dataset.
Origin Count Bb-ru mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 101 (15.9%) 4.29±2.04 6.57±3.51 4.52±1.27 7.99±2.61 4.30±2.32 10.61±2.60 6.38
U.S. 79 (12.5%) 9.76±4.95 9.99±5.25 5.51±0.71 12.01±3.53 9.04±4.33 11.97±3.75 9.71
Turkey 28 (4.4%) 5.75±2.12 9.46±1.68 3.58±1.61 11.08±8.01 10.49±2.90 16.37±2.97 9.46
Japan 68 (10.7%) 7.89±4.76 7.13±2.34 1.76±0.56 6.10±2.74 8.50±5.79 14.11±0.74 7.58
France 93 (14.7%) 7.08±3.04 6.83±2.85 5.03±1.86 7.03±2.48 8.85±4.96 10.55±3.43 7.56
U.K. 33 (5.2%) 8.42±5.82 10.03±5.01 5.54±3.74 11.89±6.16 13.17±8.89 16.82±8.31 10.98
Mexico 22 (3.5%) 4.08±0.89 2.01±0.82 2.08±0.95 4.70±1.40 5.64±3.25 4.21±1.34 3.79
India 21 (3.3%) 11.46±9.78 19.02±2.71 4.27±3.11 16.80±6.73 12.95±5.70 22.93±2.60 14.57
Germany 45 (7.1%) 7.53±5.26 7.95±3.59 4.39±2.03 9.41±2.54 11.74±4.92 11.57±4.32 8.77
China 30 (4.7%) 9.67±4.64 11.86±3.63 3.28±2.19 10.44±4.98 12.63±6.07 12.82±4.43 10.12
Iran 13 (2.1%) 10.86±6.84 11.06±5.99 15.51±4.53 22.88±6.61 10.77±8.30 23.88±6.87 15.83
Greece 18 (2.8%) 3.45±2.42 1.74±0.43 3.38±1.24 8.56±2.57 4.91±2.85 4.77±1.28 4.47
Spain 38 (6.0%) 3.35±1.51 6.51±2.61 3.37±1.37 5.27±2.52 3.88±2.07 5.61±1.84 4.67
Russia 45 (7.1%) 4.79±2.05 3.75±0.77 2.80±0.92 5.21±2.94 5.31±2.18 5.47±2.27 4.55
ALL 634 (100%) 6.85±3.14 7.76±1.83 4.28±1.11 8.84±2.67 8.19±4.04 11.52±2.82 7.91
@@
Origin Bb-ru mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 0.3319±0.04 0.3217±0.06 0.3633±0.03 0.4073±0.04 0.3774±0.02 0.3779±0.05 0.3633
U.S. 0.3796±0.04 0.3556±0.05 0.3692±0.02 0.4393±0.03 0.3938±0.05 0.4068±0.05 0.3907
Turkey 0.3199±0.06 0.3317±0.04 0.3300±0.04 0.3966±0.06 0.4012±0.02 0.4178±0.04 0.3662
Japan 0.2955±0.03 0.2796±0.04 0.2740±0.02 0.3085±0.04 0.3110±0.05 0.3319±0.01 0.3001
France 0.3438±0.05 0.3132±0.03 0.3652±0.04 0.3875±0.02 0.3954±0.04 0.3752±0.04 0.3634
U.K. 0.3647±0.08 0.3442±0.05 0.3771±0.04 0.4327±0.04 0.4149±0.08 0.4179±0.09 0.3919
Mexico 0.3150±0.01 0.3081±0.04 0.3604±0.04 0.3688±0.03 0.3909±0.02 0.3225±0.04 0.3443
India 0.3475±0.10 0.3808±0.05 0.3062±0.03 0.4375±0.06 0.3665±0.07 0.4236±0.04 0.3770
Germany 0.3500±0.08 0.3330±0.05 0.3402±0.04 0.3968±0.03 0.4053±0.04 0.3743±0.07 0.3666
China 0.3318±0.04 0.3348±0.05 0.3276±0.03 0.3741±0.07 0.3819±0.06 0.3520±0.05 0.3504
Iran 0.3252±0.06 0.3358±0.03 0.4221±0.06 0.4916±0.06 0.4418±0.04 0.4818±0.09 0.4164
Greece 0.3360±0.05 0.3015±0.04 0.3573±0.03 0.4045±0.01 0.4012±0.03 0.3365±0.05 0.3562
Spain 0.3240±0.03 0.3283±0.04 0.3625±0.04 0.3880±0.04 0.3638±0.03 0.3313±0.04 0.3497
Russia 0.3583±0.03 0.3199±0.04 0.3899±0.03 0.3732±0.04 0.3907±0.01 0.3591±0.03 0.3652
ALL 0.3395±0.04 0.3247±0.04 0.3515±0.03 0.3940±0.03 0.3824±0.03 0.3750±0.04 0.3612